whats the quantity of heat given out if a bar of brass of mass 350g is cooled from 95°c to 20°c (specific heat capacity of brass= 380j/kg–¹k–¹)
1
Expert's answer
2020-02-10T09:22:48-0500
"Q=mc(T_1-T_2)"

"Q=(0.35)(380)(95-20)=9975\ J"